2. Applying what you have learned, propose at least three possible products with mechanism from the dehydration of 2-methylcyclohexanol.

4. You will be starting with 10.0 mL of 2-methylcyclohexanol. Calculate the number of moles and mass of product you will get if the reaction is allowed to go to completion.

5. You will be analyzing both your starting material and your product(s) using IR spectroscopy. Provide a brief explanation of why this technique will show whether your reaction has been successful.
Answer #1

Q#2:

ОН CH3 - H20 CH3 H. H possibility ОН - H20 .CH3 CH3 Н H H Н CH3 possibility 3 possibility 2

Q#3:

The molecular formula (MF) of 2-methylcyclohexanol is C7H14O. The molecular weight (MW) of 2-methylcyclohexanol is 114.18 g/mol. Density of 2-methylcyclohexanol (d) = 0.93 g/mol (internet search)

Dehydration means loss of one H2O molecule from the 2-methylcyclohexanol. Thus, MF of dehydrated product = C7H12 and MW of the dehydrated product = 96.18 g/mol.

10 mL 2-methylcyclohexanol = 10mL\times0.93 g/mL= 9.3 g

The number of moles of 2-methylcyclohexanol = (9.3 g)/(114.18 g mol-1)= 8.15\times10-2 mol.

If the reaction goes to completion than 8.15\times10-2 mol dehydrated product would be obtained. Thus, the weight of the dehydrated product = number of moles \times MW = 8.15\times10-2 mol \times 96.18 g/mol = 7.84 g (answer)

Q#5:

The characteristic IR peaks for the 2-methylcyclohexanol is the OH stretching frequency at around 3300 cm-1. Whereas, the dehydrated product would exhibit an IR peak at around 1500 cm-1 characteristic for C=C stretching frequency. An IR peak at around 1500 cm-1 would be absent in the IR spectrum of 2methylcyclohexanol. Therefore, if the reaction is successful than the product from the reaction would only contain the dehydrated product. The starting material 2-methylcyclohexanol would be absent in the product. Therefore, the IR spectrum of the product would not show any peak at around 3300 cm-1 (-OH stretching frequency) but a peak at around 1500 cm-1 would be observed in the IR spectrum of the product.